Assume that a consumer’s preferences are given by u(x1, x2) = 10x11/2 *  x21/2

Currently, m = 200 and p1 = 10 and p2 = 20. Suppose now that p1 increases to p'1 = 20. What is the total effect of this price change in the optimal consumption of the two goods for the consumer, and what are the substitution and income effects?

Step 1:Solve the consumer’s problem given her preferences (described by u) and under the assumptions that m = 200, p1 = 10, p2 = 20.

The answer is (10, 5).

Step 2:

Step 2a: Find the amount of income m' that you will use in step 2b.

Step 2b: Solve the consumer’s problem given her preferences (described by u) and under the assumptions that m' = 300 (as we found in 2a) , p1 = 20, p2 = 20 (new prices).

The answer is = (7.5, 7.5).

Step 3: Solve the consumer’s problem given her preferences (described by u) and under the assumptions that m' = 200 (original income), p1 = 20, p2 = 20 (new prices).

The answer is = (5, 5).

Please help me solve this.
Answer #1

Munz ! * u(x,,H2) = 10x, 42 x 42 budget line 200 = 1000,+ 2002 at optimal point slope of u = stopte slope of budget line il Mthis n2zn, halds an = 10 . Ni 5 So, tabal effect - n-a = 5-10 = -5 Now amount of n needed to buy old bung 10 at neue price i